OpenStudy (anonymous):

Convert the radian measure to degree measure. Use the value of π found on a calculator and round answers to two decimal places. 9pi/12

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you do not need a calculator for this, very misleading question

OpenStudy (misty1212):

multiply by \(180\) and get rid of the \(\pi\) the math teachers say "divide by \(\pi\) but that means it goes away

OpenStudy (misty1212):

i.e. it is just \[\frac{9}{12}\times 180\]

OpenStudy (misty1212):

also no one writes \(\frac{9}{12}\) when you can just write \(\frac{3}{4}\) where does this question come from?

OpenStudy (anonymous):

why do you multiply by 180?

OpenStudy (anonymous):

lol my guess is FLVS

OpenStudy (misty1212):

that is a good question

OpenStudy (anonymous):

yes lol

OpenStudy (anonymous):

and i get 135

OpenStudy (misty1212):

to convert from degrees to radians, you multiply by \(\frac{180}{\pi}\) the \(\pi\)'s cancel if you have one so it is really like multiplying by \(180\)

OpenStudy (anonymous):

oh ok

OpenStudy (misty1212):

and erasing the \(\pi\) the explanation is easy

OpenStudy (misty1212):

there are \(360\) degrees in the circle measured in radians that is \(2\pi\) so to convert, multiply by \(\frac{360}{2\pi}=\frac{180}{\pi}\)

OpenStudy (anonymous):

is it really FLVS?

OpenStudy (anonymous):

yes y

OpenStudy (anonymous):

i get a huge number. i think im lost

OpenStudy (anonymous):

goes right in the "bad math" pile most of the entries are from FLVS

OpenStudy (anonymous):

no, you got the right answer

OpenStudy (anonymous):

1279100.73???

OpenStudy (anonymous):

you said you got \(135^\circ\) right?

OpenStudy (anonymous):

\[\frac{3}{4}\times 180=135\]

OpenStudy (anonymous):

no. where did that equation misty said come from with pi in it twice

OpenStudy (misty1212):

your radian measure is \[\frac{3\pi}{4}\] to convert to degrees \[\frac{3\pi}{4}\times \frac{180}{\pi}=\frac{3}{4}\times 180=135\]

OpenStudy (misty1212):

the \(\pi\)'s cancel that is why the instructions are so misleading you do not need to compute anything with the \(\pi\)